OpenStudy (anonymous):

Solve for T: 1-1/t^2-1 = 1/t^2-1

OpenStudy (anonymous):

Looks like this? \[1-\frac{1}{t^2-1}=\frac{1}{t^2-1}\]

OpenStudy (anonymous):

si

OpenStudy (anonymous):

how do i make fractions like that?

OpenStudy (anonymous):

The best thing to do is to add 1/(t^2-1) to both sides. \[1-\frac{1}{t^2-1}+\frac{1}{t^2-1}=\frac{1}{t^2-1}+\frac{1}{t^2-1}\] Notice that you have the negative of that ugly fraction and the positive of the ugly fraction on the same side. Those will add to 0. *Type frac{}{} in the equation editor. Whatever you type between the first set of {} will go on the top. For example, frac{1}{t^2-1} will look like

OpenStudy (anonymous):

You get \[1=\frac{1}{t^2-1}+\frac{1}{t^2-1}\]

OpenStudy (anonymous):

Combine like-terms: \[1=\frac{2}{t^2-1}\] Multiply both sides by t^2-1 and see what you get.

OpenStudy (anonymous):

t^2-1=2?